problem 20 I got wrong.

x + 5y =10
-2x -10y =-20
: Solve each of the following systems by addition.If a unique solution does not exist ,state whether the system is inconsistent or dependent.

Answer by fractalier(6550)

If you multiply the top equation by minus two, you go from
x + 5y =10
-2x -10y =-20
to
-2x -10y =-20
-2x -10y =-20
You can see that these are the very same line.
Thus the system is dependent and there ar an infinite number of solutions.